Meeting Summary Notes2022

From Cuawiki
Jump to: navigation, search

INTRODUCTION

  • In Experiment Table move item E12-06-114 ("DVCS days moved from Hall A") to position #3
  • included in beam time request
  • part of Phase 1, but was not part of ERR in 2019
  • uses the same equipment
  • Hall C planning process for the next few years is ongoing
  • Mark has been talking to different groups and will try to make a decision soon
  • Action item: Mark check with Patrizia Rossi to get formal approval for E12-06-114


MECHANICAL DESIGN AND STATUS

  • Detector support structure - components in house and ready for assembly this SAD
  • Cable support: method with support running underneath the SHMS platform
  • Roof block cutting planned for this SAD
  • time estimate - 2 weeks (checked with Walter)
  • hole will be covered up with additional shielding
  • NPS installation timeframes (expectation from Project estimates)
  • HB removal - 2 weeks
  • HB re-install - 3 weeks
  • NPS initial install (SHMS right) - 50 days
  • NPS removal from right side - 3 weeks
  • NPS initial install (hardware) left side - 25 days
  • NPS moved from SHMS right side to left side (looking downstream) - 1 week including 3 days of radcon cooling if immediately after a run
  • Moving the sweeper magnet in same side setup will take a work day with survey support
  • Discussion about ground during welding to protect detectors
  • welding machines will be grounded to the hall floor
  • Schedule for SAD - some concerns about physics starting in June instead of July
  • Action items:
  • check again if grounded signals/HV cables are isolated from the frame
  • Mark check with Walter and Steven L. to make sure resources are available


DETECTOR SUPPORT GROUP

  • Overview of all DSG projects completed and ongoing: HV cables and testing, ESR pre-shaping, hardware monitoring, hardware interlock system (ongoing), signal monitoring, software configuration, thermal simulation (ongoing)
  • HV cables
  • Fabricated 40 cables, need 30, so have 10 spare HV cables
  • Estimate ~30 minutes for changing HV cables (turn off all HV in crate and unplug/plug)
  • Crystal temperature depends on ambient temperature - may be of interest to have additional temperature sensors on the inside of the detector box
  • discussion about where these would go - not much space inside the box
  • cooling is outside the frame
  • Action item:
  • Make model for placement of the temperature sensors inside the box and figure out cable routing


HV DIVIDERS

  • Completed studies and now procuring components
  • hybrid approach to manufacturing
  • Procurement complete in December 2021
  • components
  • LV cables
  • PCB assembly
  • Completion goal: originally 1 May (with delays 27 May 2022)
  • one SAMTEC connector is late due to supply chain issues (current estimate is 9 March 2022)
  • removed more than half dividers (from old PCB) and plan additional recources for socket installation to mitigate the impact of the delays and speed up how many dividers can be produced per week
  • first dividers will be ready before March 27th and trying to further optimize
  • plan to test every single divider with functional tests with LED
  • Action items:
  • Further optimize divider production - all dividers will be needed at the start of the assembly and stacking will take 1 week


SOFTWARE ANALYZER DEVELOPMENT

  • Work has been completed to integrate DVCS into HCANA and infrastructure for fADC mod 10 has been implemented
  • mod10 - oscilloscope trace around trigger time
  • software can be adapted to obtain amplitude from trace
  • Discussion about DAQ setup, HV, diagnostic tools for trigger setup test, LED, etc.
  • need to define diagnostic histograms needed for cosmics tests
  • set of histograms available - check what can be used
  • Action Items:
  • need firmware development
  • need to implement event decoding in analyzer (VTP event unblocking)


CRYSTALS OVERVIEW

  • All crystals are on site and ready for assembly
  • Stacking strategy: all Crytur crystals
  • Light yield measurement done with teflon and electrical tape wrapping
  • Observe no difference in performance of crystals produced with old and new powder
  • Discussion about crystal tracking
  • goal: capture information for each module
  • web-based spreadsheet for use during assembly
  • database for longer term
  • have experience from BigCal
  • Crystal pre-shaping
  • 700 pre-shaped foils are available
  • 400 unshaped foils and tools are at JLab
  • Completion goal: have all foils ready by the end of May
  • Action Items:
  • Collect information on what columns are needed for the web-based spreadsheet
  • discuss with JLab about developing an Excel sheet with hidden columns etc.
  • find out options how to turn the web-based spreadsheet into a database
  • organize and carry out the shaping of the remaining ESR foils


CRYSTAL WRAPPING PROCEDURE

  • Estimate ~1 week for 4 people to wrap all crystals
  • 12 days for 2 people
  • Pre-assembly tasks
  • PMT+base assembly
  • PMT kapton protection
  • Assembly tools have been shipped to JLab and a step-by-step guide for the assembly is available on the Wiki
  • Storage space for crystals needs to be checked and finalized
  • see drawing from Brad's talk
  • Discussion about monitoring module sizes
  • Aim: monitor crystal size as modules are being prepared - do this row-by-row
  • do not expect problems with fit as a carbon frame is used that defines the module compartments
  • carbon compartments were made to accommodate the larger SICCAS crystal size - now that only Crytur crystals are used there should be enough space


TRAVEL AND ASSEMBLY IN CURRENT MEDCON 4 STATUS

  • For international travel (if expect support from JLab) formal approval from DOE is needed
  • timeframe: submit TA 45 days in advance of travel
  • For domestic travel (if expect support from JLab) need formal approval from director and Site Office
  • timeframe: submit forms ~2 weeks in advance - foreign nationals need to provide a CV
  • Action Item:
  • Send list of names of people planning to be on site for NPS assembly to Mark


DAQ, HV, CABLES, LAYOUT

  • Components in hand, but some items need attention, e.g., cable bundling, software development
  • Detector assembly layout in EEL
  • crystal shelving holds ~300 crystals
  • Action Items:
  • Determine what would help for crystals storage


ASSEMBLY DISCUSSION

  • the list of tasks seems complete, but needs additional detail on some items
  • add any additional items


  • Overall have a clear plan for the first 3 weeks for assembly - see the step-by-step guide and time estimates
  • DAQ/electronics test specifics
  • Aim: LED, cluter trigger tests
  • need a coordinator (Brad?) and a DAQ expert to lead the setup, as well as participants
  • HV cables for testing in EEL
  • cannot test all cables outside the hall
  • envision a special frame for these tests
  • in the hall HV goes to a patch in front of SHMS and then upstairs
  • Action Items:
  • Make a step-by-step list for the DAQ test specifics, e.g., cosmic tests, preliminary gain matching, etc.
  • Identify a DAQ expert to lead the DAQ/electronics setup (LED, cluster trigger checking)
  • Check with Jack Segal's group on cable patch panel and cable checkout
  • Prepare list of people planning to be onsite for the assembly
  • Prepare a slide/list of DAQ/electronics tests envisioned for the summer (Brad)
  • check with Brad on the HV setup to be used in EEL for testing outside the hall


PHYSICS WITH NPS

NEW IDEAS

  • coincidence parity - would use NPS as electron arm
  • 2-photon e-p with NPS as hadron arm (proton ID)
  • NPS with positron beam for A' search - 10^3 gain vs. PADME
  • Coincidence parity concept most ready
  • initial studies have been done
  • coordinate resolution of calorimeter is very good (better than lead glass) - have elastics, so don't have stringent requirements on tracking
  • need to figure out the details of the instrumentation, e.g., what should be the electron arm
  • Time estimate for Coincidence parity experiment at Q2=2-3 GeV2: About one month to get constraints on proton form factor
  • Positron experiment more difficult though in principle possible
  • front of CPS would need to be rebuilt
  • Action Item:
  • Work out coincidence parity experiment in more detail


FUTURE CALORIMETERS IN HALL C BASED ON NPS

  • Goal: large acceptance homogeneous calorimeter
  • Different possibilities to use NPS either split up into smaller calorimeters or adding crystals to the acceptance
  • Also possibility to use NPS combined with other homogeneous material like lead glass (HyCAL concept) or SciGlass (EIC goal)
  • Action Item:
  • further discuss possibilities


TCS PRESENTATIONS

TCS PROJECT

  • Polarized transverse target is on site
  • Magnet arrived in 2021
  • Hall B may have details already in GEANT4
  • Tosca models may give information on fringe fields
  • target was matched to detector setup
  • Field profile slightly modified since have 3 coils now
  • Action Items:
  • Use updated target configuration - angular acceptance should be +-25 degrees (tansverse) and +-35 degrees (longitudinal)
  • Change "collimated" to "high intensity, highly directional beam" or something along these lines

TCS IN HALL C AND IMPACTS FOR GPD

DDVCS AND J/PSI IN HALL C